OpenStudy (anonymous):

find an equation of the line L, where L is perpendicular to y=4/3x and passes through the point (3,4).

OpenStudy (anonymous):

Since it's perpendicular, the slope will be the negative recipricle of 4/3. Then you can use the forumla y-y1=m(x-x1) to find the exact equation of the line.

OpenStudy (anonymous):

i got y=-3/4x+11/4 which apparently is wrong... is this correct?